OpenStudy (juan1857):

Write the expression as a single natural logarithm. 3ln3+3lnc

geerky42 (geerky42):

\[\log(a)+\log(b)=\log(ab)\]So we have \(3\ln(3)+3\ln(c)=3[\ln(3)+\ln(c)]=\boxed{3\ln(3c)}\)

OpenStudy (juan1857):

for some reason MathWay Calculator says ln(27c^3)

jhonyy9 (jhonyy9):

yes this is bc. 3ln(3c) = ln(3c)^3 = ln(3^3 *c^3) = ln(27c^3) hope this is really clearly understandably now

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (juan1857):

let me try to process this

OpenStudy (juan1857):

oh ok so which one is the true answer?

jhonyy9 (jhonyy9):

every two bc. 3ln(3c) = ln(27c^3)

OpenStudy (juan1857):

ok. Thank You

jhonyy9 (jhonyy9):

but this is the end ln(27c^3) bc. this not can simplifie it again like the 3ln3c ok. ?

OpenStudy (juan1857):

ok